Calculate the cash receipts expected for the month of March. Collection patterns indicate 75% of the month's sales will be collected in the month of the sale, 20% in the following month, and 2 % two months after the sale. Sales forecasts are $250,000 in January, $325,000 in February, and $285,000 in March.
Answer #1

Answer: Cash receipts fr the month of march is $283,750

January sales = 250,000 * 75% = $187,5000 which is collected in the same month

250,000 * 20% = $50,000 which is collected in the following month i.e February month

250,000 * 2% = $5000 which is collected from two moths of sales i.e march month

February sales = 325,000 * 75% = $243,750 in the same month

= 325,000 * 20% = $65,000 in the month of march

March sales = 285,000 * 75% = $213,750 in the same month.

Total of bold area is march month cash collection i.e 5000 + 65,000 + 213,750 =$283,750